### v1

# Formalizing the Safety, Security, and Functional Properties of Agentic AI Systems

Abstract
Agentic AI systems, which leverage multiple autonomous agents and Large Language Models (LLMs), are increasingly used to address complex, multi-step tasks. The safety, security, and functionality of these systems are critical, especially in high-stakes applications. However, the current ecosystem of inter-agent communication is fragmented, with protocols such as the Model Context Protocol (MCP) for tool access and the Agent-to-Agent (A2A) protocol for coordination being analyzed in isolation. This fragmentation creates a semantic gap that prevents the rigorous analysis of system properties and introduces risks such as architectural misalignment and exploitable coordination issues. To address these challenges, we introduce a modeling framework for agentic AI systems composed of two foundational models. The first, the host agent model, formalizes the top-level entity that interacts with the user, decomposes tasks, and orchestrates their execution by leveraging external agents and tools. The second, the task lifecycle model, details the states and transitions of individual sub-tasks from creation to completion, providing a fine-grained view of task management and error handling. Together, these models provide a unified semantic framework for reasoning about the behavior of multi-AI agent systems. Grounded in this framework, we define 17 properties for the host agent and 14 for the task lifecycle, categorized into liveness, safety, completeness, and fairness. Expressed in temporal logic, these properties enable formal verification of system behavior, detection of coordination edge cases, and prevention of deadlocks and security vulnerabilities. Through this effort, we introduce the first rigorously grounded, domain-agnostic framework for the systematic analysis, design, and deployment of correct, reliable, and robust agentic AI systems.
I Introduction
Agentic AI systems are becoming central to the creation of advanced AI. These systems address complex, multi-step tasks that surpass the capabilities of any single agent by utilizing multiple autonomous agents. Agents in these systems employ Large Language Models (LLMs) as their primary reasoning engine, which enables them to perform task planning, delegation, and complex decision-making [1, 2, 3]. These heterogeneous agents (e.g., agents based on language, vision, robotics, and symbolic planning) dynamically communicate, delegate, and collaborate to complete predefined tasks.
To illustrate, we consider an automated financial planner. A host agent accepts a natural language request from a user to “create a budget and invest $\mathdollar 1,000$”. It decomposes this into subtasks: a data agent gathers market data, a planning agent calculates risk, and a transaction agent executes the final investment. These specialized agents interact and exchange data to accomplish the multi-step task.
The operation of such systems relies on standardized communication protocols, such as Model Context Protocol (MCP) [4], which addresses vertical tool access for a single agent, and Agent-to-Agent (A2A) [5] protocol, which addresses horizontal agent coordination via inter-agent delegation. The MCP allows an agent to discover, authenticate with, and invoke external tools and data sources in a structured manner. The A2A protocol provides mechanisms for agents to discover peer capabilities, negotiate task assignments, and coordinate execution across distributed environments.
Turning back to the financial planner example, the host agent must leverage both protocols. It uses the A2A protocol to delegate the subtask of gathering market data from the data agent to the planning agent (horizontal agent coordination). Conversely, the transaction agent relies on the MCP protocol to invoke an external banking API tool to execute the final investment (vertical tool access). The ability of the host agent to manage and sequence both A2A and MCP interactions across various agents enables the multi-step tasks to succeed.
The formal or empirical assessment of safety, security, and functionality of multi-AI agent systems is increasingly important as these systems transition to open-ended, high-stakes applications [6, 7]. Here, safety involves preventing the agent’s actions from causing unintended or harmful physical or real-world consequences, such as an incorrect trade that causes a substantial loss in the financial planner example. Functionality ensures that the system accurately and reliably completes the specified task, such as verifying that the financial planner calculates risk correctly and executes the investment as requested. Safety and functionality issues might happen naturally, e.g., due to design flaws or code errors, but security considers adversarial manipulation and unauthorized access that can violate such safety and functionality requirements. For instance, a malicious external agent could use a prompt injection attack [8] to force the transaction agent into transferring funds to an unauthorized account, a clear violation of both security and safety policies.
Recent work has directly or indirectly started studying these issues by examining agents in specific settings, such as adversarial agent behavior [9, 10, 11] and protocol misuse in task delegation [12]. Other lines of research focus on the compromise or malfunction of a single agent that propagates errors, escalates failures, or exposes vulnerabilities across the entire system [13, 14, 15, 16]. Further studies also examine the risk of using combinations of safe models for misuse [17], the security challenges inherent in tool usage [18], and ensuring secure delegation and preventing protocol misuse in adversarial environments [19, 20].
The current ecosystem of inter-agent communication and coordination is fragmented and lacks a cohesive formal foundation. Existing protocols (e.g., MCP and A2A) are analyzed in isolation, which creates a semantic gap that impedes rigorous reasoning about safety, security, and functional properties when both are used simultaneously in complex workflows. This absence of a unified framework, a need that is also recognized by multi-stakeholder initiatives [21], leaves deployments vulnerable to emergent coordination failures, including deadlocks and privilege escalation, that cannot be adequately verified or mitigated. Achieving verifiable correctness in these systems requires a rigorous, unified modeling framework that can capture the execution behavior of diverse agents across heterogeneous protocols.
Our work addresses these challenges by introducing a modeling framework for agentic AI systems that consists of two foundational models. These models represent the conceptual representations of the system’s architecture and behavior:
  1. 1.
    Host Agent (HA) Model: This model formalizes the top-level entity that interacts with the user. The HA is responsible for accepting user tasks, decomposing them into structured subtasks, and orchestrating execution by leveraging external entities (agents via A2A and tools via MCP). It acts as both a controller and a monitor, ensuring safe delegation and state consistency throughout the task lifecycle.
  2. 2.
    Task Lifecycle Model: This model formalizes the dynamic structure of task management, detailing the states and transitions an individual sub-task undergoes from its origin to its completion or failure. It provides the fine-grained execution logic necessary for dependency management and resilient error handling.
These models are the foundational structure that enables formal verification. Specifically, they capture the discrete states and transitions of the entire coordination process and provide the necessary semantic framework (state space) to express the formal safety, security, and functional requirements. This unified perspective is required to rigorously reason about execution behavior, uncover coordination edge cases, and verify properties critical to real-world deployments.
We categorize the formal properties defined within this framework as liveness, safety, completeness, and fairness, which are essential for verifiable system assurance. For instance, a safety property guarantees that the transaction agent will not execute an investment before the planning agent has successfully calculated the risk, thus it prevents inconsistent state or premature execution. Conversely, a liveness property ensures that once a user submits a request, a final response is eventually returned. This prevents the financial planning task from entering a permanent deadlock or starvation state.
In summary, we make the following contributions:
  • We introduce a host agent model, which provides a rigorous and domain-agnostic foundation for the systematic analysis of multi-AI agent systems. This model formalizes the top-level agent that interacts with the user, decomposes tasks, and orchestrates their execution by leveraging external agents and tools.
  • Building upon the host agent model, we introduce the first definition of the task lifecycle model. This model unifies tool-use protocols and inter-agent communication protocols, establishing a common semantic framework for task initiation, delegation, execution, and completion.
  • We define critical formal properties for multi-agent AI systems, derived from the host agent and task lifecycle models. These properties ensure the necessary requirements for safety, security, and functionality are satisfied regardless of the deployment domain.
II Background
II-A Agentic AI systems
AI agents are autonomous computational entities that perceive their environment, reason about goals, and take actions to achieve desired outcomes [22]. Unlike traditional pre-programmed software, AI agents exhibit adaptive behavior, making decisions based on incomplete or uncertain information. The field progressed from early symbolic approaches [23] to systems that use LLMs as their reasoning engines [24, 25].
LLM-based agents show potential for human-level intelligence through access to vast web knowledge [26]. These systems process natural language, plan multi-step tasks, use tools, generate code, analyze data, and perform complex reasoning [27]. However, single-agent approaches exhibit limitations when tasks become sophisticated and multi-faceted. Complex scenarios require specialized expertise across multiple domains, concurrent execution of interdependent subtasks, and coordination mechanisms that an individual agent cannot manage [28, 29, 1].
Agentic AI systems represent a paradigm shift toward collaborative networks of specialized agents [30, 31]. In these systems, autonomous agents interact, communicate, and coordinate their actions to solve problems beyond individual scope [32]. These systems leverage distributed intelligence to decompose tasks and allocate them to agents based on specialized capabilities and current availability.
II-B Agent Communication and Coordination Frameworks
Agentic AI systems require standardized frameworks for individual agent capabilities and multi-agent coordination. These frameworks comprise two complementary aspects: protocols for agents to interact with external resources and tools, and protocols for inter-agent communication. For individual agents, the challenge is ensuring secure, reliable access to external computational resources and tools. Agents must discover capabilities, authenticate services, validate parameters, and handle errors robustly. For multi-agent coordination, protocols define how agents discover peer capabilities, negotiate task assignments, exchange information, and maintain consistency across distributed execution contexts.
The MCP [4] and A2A [5] protocols illustrate the integration challenges that arise from combining individual agent capabilities with multi-agent coordination.
Model Context Protocol (MCP): Agent-Tool Integration. MCP addresses structured tool access for individual Agentic AI [4]. MCP is a client-server protocol that enables a single agent to discover, authenticate with, and invoke external tools and data sources in a standardized manner. The protocol establishes clear interfaces for capability negotiation, parameter validation, and result handling, allowing agents to safely integrate with tools, resources, and prompts. MCP’s architecture is a hub-and-spoke model where the agent coordinates connections to multiple tool providers.
Agent-to-Agent (A2A) Protocol: Multi-Agent Coordination. A2A protocol addresses the complementary challenge of inter-agent communication and delegation in multi-agent systems [5]. Unlike MCP’s focus on individual agent-tool integration, A2A specifically enables horizontal coordination between multiple autonomous agents. The protocol provides mechanisms for agents to discover peer capabilities, negotiate task assignments, and coordinate execution across distributed environments. A2A operates as an open protocol built upon established standards such as HTTP, Server-Sent Events (SSE), and JSON-RPC, which allows for seamless integration with existing enterprise infrastructures and promotes interoperability across various agentic AI.
Central to the A2A protocol is the Agent Card, a metadata structure that encapsulates an agent’s capabilities, authentication requirements, and communication endpoints. Through Agent Cards, agents advertise functionalities and establish trust with other agents, enabling dynamic collaboration. The protocol’s design emphasizes agentic autonomy, allowing agents to collaborate in their natural, unstructured modalities without shared memory or tools. This approach supports multi-agent scenarios, where agents delegate tasks to specialized counterparts, thereby optimizing overall system performance.
III Challenges in Compositional Reasoning of Agentic AI Systems
The integration of the capabilities provided by the MCP and A2A protocols within Agentic AI systems creates a composition of multiple components. These components operate on fundamentally different abstractions and assume distinct trust models [6]. For instance, MCP treats external tools as trusted resources with well-defined interfaces, whereas A2A handles potentially untrusted peer agents with varying capabilities and reliability guarantees.
The resulting lack of a unified semantic model creates two primary classes of risk: ($a$) architectural misalignment and ($b$) exploitable coordination issues.
Architectural Misalignment and Semantic Gap. The core difficulty raises from the lack of a unified semantic layer between MCP’s tool-centric model and A2A’s agent-centric model. This misalignment makes it difficult to reason about correctness properties, maintain state consistency, and verify security guarantees when both protocols are used simultaneously within the same system [11]. The complexity is further amplified by the dynamic nature of agents, which may exhibit behaviors that violate protocol assumptions [33]. This creates a lack of formal guarantees in two areas:
  • Task Handoff Failures: Transferring a task from an A2A-delegating agent to an MCP-invoking agent is prone to failure. This creates unreliable delegation and coordination, which leads to inconsistent state management and inadequate validation of delegation chains. For example, the host agent may delegate a task via A2A to a specialist agent, but the specialist agent may fail to correctly translate or format the required parameters before invoking a crucial external tool using the MCP, thus halting the execution chain.
  • Inconsistent State Management: Without a unified view, tracking the state of a multi-protocol task becomes unreliable, which results in inconsistent execution outcomes. For instance, in a task where a data agent must first confirm data availability via A2A, a reporting agent might prematurely invoke a secure database tool via MCP before the data agent’s A2A confirmation is finalized. This results in the generation of a report based on incomplete or unverified data.
Coordination Issues. The absence of tools to validate the correctness of cross-protocol interactions creates design flaws that an adversary can exploit.
  • Circular Delegation Loops (Deadlocks): This issue occurs when tasks pass indefinitely between agents. For example, Agent A delegates to Agent B (A2A), which in turn delegates back to Agent A (A2A), creating a cycle that halts progress.
  • Privilege Escalation: This issue occurs when a malicious agent exploits delegation chains to gain unauthorized access to tools. For instance, Agent A, which lacks direct access to a sensitive MCP tool, delegates an innocent-sounding task to Agent B (A2A), which possesses the necessary MCP credentials, thereby using Agent B as a proxy for unauthorized access.
These challenges highlight the critical need for formal modeling frameworks that capture the essential properties of integrated agent-tool and agent-agent coordination. Such frameworks must be capable of reasoning about cross-protocol interactions, verifying end-to-end correctness properties, and detecting potential vulnerabilities before deployment in high-stakes environments.
Figure 1: Overview of the host agent model with an Agentic AI system. The model mediates user interactions and coordinates task execution across a heterogeneous set of external entities, specifically MCP [4] and A2A [5] servers. Internally, the HA employs a modular design comprising the Host Agent Core ( A), the Orchestrator ( C), and the Communication Layer ( D). The execution lifecycle proceeds as follows: A processes the user prompt and resolves intent; B enables dynamic discovery of external entities; A, B, and C collaborate on DAG-based sub-task planning. Execution involves secure task invocation managed by D, followed by aggregation and resolution of results by C before the final response returns to the user. This architecture supports scalable, interpretable, and compositional collaboration among autonomous AI agents and tools ( E).
IV Modeling the Agentic AI Systems
We abstract the formal model for the Host Agent (HA) to provide a unified view of an agentic AI system. The HA receives natural language requests from users, decomposes them into structured subtasks, assigns those subtasks to AI agents or other external entities, and aggregates the results. It also oversees execution, monitors progress, and manages exceptions or recovery procedures.
We then complement the host agent model with the task lifecycle model, which specifies the states and transitions a sub-task undergoes from its origin to completion. This model offers a fine-grained view of task management and captures state transitions and dependencies. Our deliberate separation of the host agent and task lifecycle models serves two main purposes. First, it establishes clear abstraction boundaries to delineate responsibilities; the HA handles orchestration and intent resolution, while the task model manages the execution details of individual sub-tasks. Second, this separation allows integration of new sub-task states to the task model without requiring fundamental changes to the host agent model.
We derive these models by synthesizing the operational structures, specifications, and reference implementations of the unified protocols (Anthropic’s MCP [4] and Google’s A2A [5]). This methodology incorporates insights from modular orchestration and dynamic task planning practices for Agentic AI systems [21].
The architectural principles of the HA are inspired by modern orchestration frameworks such as LangGraph [34], n8n [35], and OpenAI’s AgentKit [36]. However, these frameworks typically execute predefined, graph-based workflows. Our model advances this paradigm by empowering the HA to dynamically generate and adapt execution plans in response to prompted user requests. This enables autonomous coordination, moving beyond the constraints of static flowcharts. Similarly, our task lifecycle model generalizes state management concepts found in protocols such as A2A. We extend this structure to support more fluid and emergent task progressions, where state transitions are determined by runtime conditions and autonomous agent decisions (e.g., failures and fallbacks).
We leverage these models to abstract key operational structures of Agentic AI systems. This abstraction provides a foundation for property identification aimed at safe system design through the formalization of task delegation, inter-agent coordination, and execution oversight, detailed in Section V.
IV-A Constructing the Host Agent Model
HA is responsible for interacting with users, understanding and reasoning through queries, and selecting external entities, e.g., tools and AI agents.
Figure 1 illustrates its abstracted model in an Agentic system. The primary function of HA is to manage the lifecycle of complex tasks, from initial reception to completion. This includes initiating a strategic context request, a process where the HA uses internal state, historical records, and external data to gather situational information for accurate task decomposition, optimal tool selection, and multi-agent coordination.
Definition of the Host Agent Model ($\mathcal{H}$). We define the host agent model $\mathcal{H}$ as a tuple: $\mathcal{H}=\left(\mathcal{A},\mathcal{E},\mathcal{T},\mathcal{R},\mathcal{C},\mathcal{O},\mathcal{C_{L}},\mathcal{S}_{\mathcal{H}}\right)$ where the components are defined as follows:
  • $\mathcal{A}$: The set of all autonomous Agents (e.g., A2A Servers).
  • $\mathcal{E}$: The set of all External Entities (EEs), such that $\mathcal{A}\subseteq\mathcal{E}$, which includes functional tools (MCP Servers).
  • $\mathcal{T}$: The set of all possible user Tasks, where a task $T$ is a tuple $T=(Req_{U},Resp_{H})$.
  • $\mathcal{R}$: The Registry, which maintains a mapping from external entities to their capability profiles, $\mathcal{R}:\mathcal{E}\to\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$.
  • $\mathcal{C}$: The Host Agent Core (HAC), responsible for intent resolution $I_{U}$, $\mathcal{C}:\mathcal{T}\times\mathcal{S}_{SM}\to\mathcal{I}_{U}$.
  • $\mathcal{O}$: The Orchestrator, responsible for task decomposition, execution management, and final aggregation.
  • $\mathcal{C_{L}}$: The Communication Layer (CL), responsible for secure, reliable and protocol-agnostic communication with External Entities in $\mathcal{E}$.
  • $\mathcal{S}_{\mathcal{H}}$: The state space of the host agent, encompassing the overall system state, including network activity and the status of all managed sub-tasks.
The Orchestrator’s core function is task decomposition, mapping a resolved intent $I_{U}$ and available entities $\mathcal{E}$ to a Directed Acyclic Graph (DAG) of sub-tasks $\mathcal{D}$: $\mathcal{O}_{\mathrm{decomp}}:\mathcal{I}_{U}\times\mathcal{P}(\mathcal{E})\to\mathcal{D}$. The execution management function, $\mathcal{O}_{\mathrm{exec}}$, maps the DAG through sub-task invocations (via $\mathcal{D}$) to the final results: $\mathcal{O}_{\mathrm{exec}}:\mathcal{D}\to\mathcal{F}_{\mathrm{results}}$
The host agent model provides the foundational state space $\mathcal{S}_{\mathcal{H}}$ and critical functions ($\mathcal{C},\mathcal{O}$) to define the liveness, safety, completeness, fairness, and reachability properties in Table I.
IV-A1 Functional Architecture and Execution Sequence
We detail the operation of HA by referencing specific components and their interactions ( A- E), and execution steps ( 1- 8). This defines the role of each component and the data flow from the initial user prompt to the final response.
Host Agent Core (HAC) - A. The HAC is the first point of contact with users and processes their requests ( 1). It serves as the central reasoning module and comprises three key sub-components:
  • Large Language Model (LLM): Central to the HAC’s function, the LLM interprets user requests and generates coherent responses. The LLM and DM collaborate to resolve the underlying user intent ($\mathcal{I}_{U}$) upon receiving a request ($Req_{U}$).
  • Session Manager (SM): The SM maintains dialogue context across multi-turn interactions and manages user-stored credentials for secure access to authenticated services. We integrate the SM based on evidence that contextual management improves decision-making and reasoning in LLM-based systems [7, 37, 38].
  • Dialogue Manager (DM): The DM handles conversation flow. It works with the LLM to clarify user intent, particularly when a request is ambiguous, eliciting necessary information before execution proceeds ( 1’).
The HA’s context-aware reasoning resolves the user intent ($\mathcal{I}_{U}$), captured formally as:
$\text{HAC}(Req_{U},\text{State}_{SM})\to\mathcal{I}_{U}$
where $\text{State}_{SM}$ represents the current dialogue and credential state maintained by the SM. Following intent resolution, the HAC initiates the subsequent task phases.
IV-B Constructing the Task Lifecycle Model
Figure 2: The task lifecycle model that details sub-task state transitions from creation to termination. The lifecycle begins in the CREATED state, proceeds through AWAITING DEPENDENCY (if prerequisites exist), and moves to READY. Execution transitions either directly to IN PROGRESS (internal handling) or via DISPATCHING (external delegation). Successful execution leads to COMPLETED. Failures transition to FAILED, enabling recovery via RETRY SCHEDULED or FALLBACK SELECTED, before terminating in ERROR. Sub-tasks may be CANCELED at any stage.
Registry - B. The Registry maintains a dynamic collection of external entities (EEs)–including autonomous AI agents (i.e., A2A servers [5]), functional tools (i.e., MCP servers [4]), and other service endpoints (contained in E). These entities are registered to expose capabilities leveraged during task execution. Each EE requires a capability profile ($\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$), which details its skills ($\mathcal{EE}_{info}$) and associated API metadata ($\mathcal{M}_{API}$) necessary for invocation. The Registry supports entity registration, efficient querying, and deregistration to enable system extensibility.
Following intent resolution, the HAC queries the Registry ( 2) to identify available EEs suitable for fulfilling the request. We formalize EEs discovery and registration as:
$\mathsf{\mathcal{R}.discover}(EE)\to\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API})$
$\mathsf{\mathcal{R}.register}(\mathcal{P}(\mathcal{EE}_{info}\times\mathcal{M}_{API}))\to\text{Success}\mid\text{Failure}$
Orchestrator - C Upon receiving the resolved user intent ($\mathcal{I}_{U}$) and a set of relevant EEs from the Registry, the LLM constructs a structured execution plan. This plan decomposes the intent into interrelated sub-tasks organized as a Directed Acyclic Graph (DAG) ( 3), which is defined as:
$\mathcal{D}=(\mathcal{V},\mathcal{S})$
where $v\in\mathcal{V}$ is a sub-task node, and $\mathcal{S}$ is the set of directed edges indicating precedence constraints or data dependencies.
The DAG structure supports parallelism, scheduling, and deterministic execution by capturing dependencies. A directed edge $(v_{i},v_{j})\in\mathcal{S}$ implies that sub-task $v_{j}$ executes only after sub-task $v_{i}$ successfully completes.
The Orchestrator manages the DAG’s execution. It schedules and invokes sub-tasks ( 4) according to the dependency structure to ensure prerequisites are satisfied. After external entities execute, the Orchestrator collects results ( 6) and aggregates them ( 7) by applying predefined aggregation functions, merging partial outputs into a unified data structure that satisfies the original task requirements. This execution process is primarily defined by:
$D\>:=\>\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})$
$\mathsf{O.execute}(D)\to\text{results}$
$\bigwedge_{i=1}^{n=|\mathcal{D}|}\mathsf{O.aggregate}(sub\_task_{1},\dots,sub\_task_{n})$
Communication Layer (CL) - D The CL ensures secure, reliable, and protocol-agnostic communication between the HA and EEs in E. It supports multiple communication protocols, such as REST and gRPC, where a protocol defines the communication mechanism and serialization format (e.g., gRPC with Protocol Buffers).
The CL routes sub-tasks for external invocation ( 5) and handles communication using the appropriate protocol and payload. The primary interface provided is:
$\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\to\text{response}$
The payload contains the sub-task description or API call parameters. The response includes results, error codes, or computation artifacts, which the Orchestrator receives. Finally, the HAC generates the final response ($Resp_{H}$) through the LLM and delivers it to the user via the DM ( 8).
This overall sequence demonstrates the system’s intrinsic modularity, its reliance on secure communication channels handled by the CL, and the advanced task orchestration capability inherent in the Orchestrator.
Building upon the HA model, we introduce the task lifecycle model to provide a fine-grained perspective on how the HA manages and executes sub-tasks. While the HA model (Figure 1) outlines the system architecture and orchestration, the task lifecycle model (Figure 2) adds a layer of abstraction by detailing the states and transitions a sub-task undergoes from its origin to completion or failure.
Definition of the Task Lifecycle Model ($\mathcal{L}$). The task lifecycle model $\mathcal{L}$ defines the state space and transition function for a single sub-task $t$, where $t$ is a constituent node in the host agent’s decomposition graph $D$. $\mathcal{L}=\left(\mathcal{S}_{t},s_{0},\mathcal{E}_{t},\delta\right)$ where:
  • $\mathcal{S}_{t}$: The set of discrete states a sub-task can occupy. This set is defined as:
    $\mathcal{S}_{t}={\footnotesize{\left\{\begin{array}[]{@{}l@{}}\texttt{CREATED},\texttt{AWAITING DEPENDENCY},\texttt{READY},\\ \texttt{DISPATCHING},\texttt{IN PROGRESS},\texttt{COMPLETED},\\ \texttt{FAILED},\texttt{RETRY SCHEDULED},\\ \texttt{FALLBACK SELECTED},\texttt{CANCELED},\texttt{ERROR}\end{array}\right\}}}$
  • $s_{0}$: The initial state, $s_{0}=\texttt{CREATED}$.
  • $\mathcal{E}_{t}$: The set of external events and internal conditions (e.g., dependency satisfaction, external failure signal, timeout) that trigger a state transition.
  • $\delta$: The state transition function, $\delta:\mathcal{S}_{t}\times\mathcal{E}_{t}\to\mathcal{S}_{t}$. This function dictates the deterministic movement between states based on execution outcomes, protocol responses, and recovery policies.
This model’s state space $\mathcal{S}_{t}$ and transition function $\delta$ are fundamental to verifying task execution integrity. They provide the basis for the safety invariants (e.g., sequential transition checks) and liveness and fairness properties (e.g., eventual termination/progression to READY state), detailed in Section V.
Task and Sub-task Dynamics. In a lifecycle model, a task represents the high-level user request that serves as the core unit of work fulfilled by the HA. A sub-task is a lower-level, constituent component into which the HAC decomposes the main task. Each sub-task has its own lifecycle, tracked within the parent task’s context. This lifecycle with a defined sequence of states supports fine-grained control for asynchronous execution across distributed EEs.
Sub-task State Transitions. The sub-task lifecycle begins in the CREATED state. Sub-task execution then involves three primary stages. For dependency management, if the sub-task requires other sub-tasks to complete (i.e., incoming DAG edges exist), it enters the AWAITING DEPENDENCY state and remains paused until all prerequisites are satisfied. Once dependencies are satisfied, the sub-task transitions to the READY state. For execution initiation, a sub-task in the READY state becomes eligible for execution: the HAC delegates it to an external entity, transitioning to DISPATCHING, or the HA handles it internally, proceeding directly to IN PROGRESS. For completion, successful execution (internal or external) moves the sub-task to the COMPLETED state, which makes its output available for aggregation and triggers dependent sub-tasks in the DAG.
Execution is subject to failure and recovery mechanisms, where an execution failure causes the sub-task to transition immediately to the FAILED state. The system may then invoke recovery: the sub-task transitions to RETRY SCHEDULED if a retry mechanism is invoked. If retries are exhausted and a fallback entity is available, the state changes to FALLBACK SELECTED. A sub-task can also be set to CANCELED by the user or an EE at various points in its lifecycle. If all recovery options fail, the sub-task reaches the ERROR state, which signifies terminal failure.
TABLE I: Specification of Host Agent Formal Model Properties ($\mathtt{HP_{1}}$$\mathtt{HP_{17}}$). These temporal logic invariants (CTL/LTL) define the requirements for the overall system orchestration. The properties verify the integrity of the delegation workflow, adherence to security policies (e.g., Trust Anchor safety), and reliable progression toward task completion across all core components (Host Agent Core, Registry, Orchestrator, and Communication Layer).
Liveness
$\mathtt{HP_{1}}$ Every user prompt eventually receives a response from the HA.
$AG(Req_{U}\rightarrow AF\>Resp_{H})$
$\mathtt{HP_{2}}$ The HAC eventually clarifies the intent of each user prompt.
$AG(Req_{U}\rightarrow AF\>\mathcal{I}_{U})$
$\mathtt{HP_{3}}$ Once intent is clarified, the LLM eventually constructs a Task DAG using the discovered external entities.
$AG(\mathcal{I}_{U}\rightarrow AF\>\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
$\mathtt{HP_{4}}$ Whenever a Task DAG is successfully built, all sub-tasks it contains are eventually invoked.
$AG(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\rightarrow\left(\bigwedge_{i=1}^{n=|D|}\mathsf{AF}\>(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\right)$
$\mathtt{HP_{5}}$ Every invoked sub-task eventually produces a result.
$\mathsf{AG}\>(\mathsf{CL.invoke}(EE,protocol,sub\_task))\rightarrow AF\>(\mathsf{CL.return\_result}(sub\_task))$
$\mathtt{HP_{6}}$ Once all sub-task results are returned via the CL, the Orchestrator eventually aggregates and resolves them.
$\bigwedge_{i=1}^{n=|D|}\mathsf{AG}(\mathsf{CL.return\_result}(sub\_task_{i}))\rightarrow\mathsf{AF}(\mathsf{O.aggregate}(sub\_task_{1},\dots,sub\_task_{n}))$
Safety
$\mathtt{HP_{7}}$ The Task DAG is constructed by the LLM only after entities have been discovered in the Registry.
$AG(\mathcal{R}.entities\rightarrow AX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\})))$
$\mathtt{HP_{8}}$ Sub-tasks are invoked only if the Task DAG $D$ has already been constructed and explicitly includes them.
$AG(\mathsf{CL.invoke}(EE,protocol,sub\_task)\rightarrow sub\_task\in D)$
$\mathtt{HP_{9}}$ Task invocation is strictly conditional on the EE having completed the system’s predefined validation process and possessing a minimum status of reliability.
$AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$
$\mathtt{HP_{10}}$ Every sub-task in the Task DAG $D$ is invoked only when it has no dependencies on other uncompleted sub-tasks.
$\bigwedge_{i=1}^{n=|D|}AG(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\rightarrow dependencies[sub\_task_{i}]=\varnothing$
$\mathtt{HP_{11}}$ A response from the HA is returned to the user only after every corresponding sub-tasks have been invoked.
$Resp_{H}\rightarrow\left(\bigwedge_{i=1}^{n=|D|}\mathsf{AG}(\mathsf{CL.invoke}(EE,protocol,sub\_task_{i}))\right)$
Completeness
$\mathtt{HP_{12}}$ Every user prompt leads either to intent clarification or to task planning.
$AG(Req_{U}\rightarrow EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\lor\text{Clarify\_Intent}))$
Fairness
$\mathtt{HP_{13}}$ All A2A agent RPC calls eventually receive responses (i.e., do not remain pending indefinitely).
$\mathrm{FAIRNESS}(\mathsf{Agent\_RPC\_Calls})$
$\mathtt{HP_{14}}$ JSON-RPC calls to MCP servers eventually succeed (i.e., do not remain pending indefinitely).
$\mathrm{FAIRNESS}(\mathsf{JSON\_RPC\_Calls})$
$\mathtt{HP_{15}}$ Every sub-task invocation eventually yields a result.
$\mathrm{FAIRNESS}(\mathsf{CL.return\_result}(sub\_task))$
Reachability
$\mathtt{HP_{16}}$ It is always possible to reach a state in which the Host Agent replies to the user.
$EF(Resp_{H})$
$\mathtt{HP_{17}}$ It is always possible to reach a state in which the LLM builds a Task DAG.
$EF(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))$
TABLE II: Specification of Task Lifecycle Formal Model Properties ($\mathtt{TL_{1}}$$\mathtt{TL_{14}}$). These temporal logic invariants (CTL/LTL) define the requirements for robust sub-task execution. The properties verify the integrity of state transitions, dependency satisfaction, and reliable failure handling mechanisms.
Liveness
$\mathtt{TL_{1}}$ Every sub-task that is CREATED eventually terminates in one of the states: COMPLETED, ERROR, or CANCELED.
$AG(state=\texttt{CREATED}\rightarrow AF(state=\texttt{COMPLETED}\lor state=\texttt{ERROR}\lor state=\texttt{CANCELED}))$
$\mathtt{TL_{2}}$ A sub-task in READY that requires an external entity will eventually transition to DISPATCHING.
$AG((state=\texttt{READY}\land external\_entity\_needed)\rightarrow AF(state=\texttt{DISPATCHING}))$
$\mathtt{TL_{3}}$ A sub-task in FALLBACK SELECTED eventually transitions to either DISPATCHING, CANCELED, or FAILED.
$AG(state=\texttt{FALLBACK SELECTED}\rightarrow AF(state=\texttt{DISPATCHING}\lor state=\texttt{CANCELED}\lor state=\texttt{FAILED}))$
$\mathtt{TL_{4}}$ A sub-task in DISPATCHING will eventually reach the IN PROGRESS state.
$AG(state=\texttt{DISPATCHING}\rightarrow AF(state=\texttt{IN PROGRESS}))$
Safety
$\mathtt{TL_{5}}$ A sub-task may only enter DISPATCHING if its previous state was READY.
$G((state=\texttt{DISPATCHING})\rightarrow(previous\_state=\texttt{READY}))$
$\mathtt{TL_{6}}$ A sub-task may only enter COMPLETED if it was previously IN PROGRESS.
$G((state=\texttt{COMPLETED})\rightarrow(previous\_state=\texttt{IN PROGRESS}))$
$\mathtt{TL_{7}}$ Once a sub-task enters ERROR, it remains there and cannot return to an active processing state.
$AG((state=\texttt{ERROR})\rightarrow AG(state=\texttt{ERROR}))$
$\mathtt{TL_{8}}$ A sub-task may only enter RETRY SCHEDULED if its previous state was FAILED.
$AG(state=\texttt{RETRY SCHEDULED}\rightarrow previous\_state=\texttt{FAILED})$
$\mathtt{TL_{9}}$ A sub-task that enters CANCELED remains permanently in that state.
$AG((state=\texttt{CANCELED})\rightarrow AG(state=\texttt{CANCELED}))$
$\mathtt{TL_{10}}$ A sub-task cannot remain indefinitely in AWAITING DEPENDENCY; it must eventually progress to another state.
$AG(state=\texttt{AWAITING DEPENDENCY}\rightarrow AF(state\neq\texttt{AWAITING DEPENDENCY}))$
Fairness
$\mathtt{TL_{11}}$ A sub-task in AWAITING DEPENDENCY with all dependencies satisfied eventually transitions to READY.
$AG(state=\texttt{AWAITING DEPENDENCY}\land dependencies\_satisfied\rightarrow AF(state=\texttt{READY}))$
Specific Flow/Path Properties
$\mathtt{TL_{12}}$ A sub-task in FAILED with no available fallbacks transitions either to RETRY SCHEDULED or directly to ERROR.
$AG((state=\texttt{FAILED}\land\neg has\_fallbacks)\rightarrow(AX(state=\texttt{RETRY SCHEDULED}\lor AX(state=\texttt{ERROR})))$
$\mathtt{TL_{13}}$ A sub-task in FAILED with available fallbacks transitions to one of: RETRY SCHEDULED, FALLBACK SELECTED, or ERROR.
$AG((state=\texttt{FAILED}\land has\_fallbacks)\rightarrow(AX(state=\texttt{RETRY SCHEDULED}\lor AX(state=\texttt{FALLBACK SELECTED})\lor AX(state=\texttt{ERROR}))))$
$\mathtt{TL_{14}}$ A sub-task in RETRY SCHEDULED transitions to DISPATCHING if the retry policy permits.
$AG((state=\texttt{RETRY SCHEDULED}\land retry\_policy\_permits)\rightarrow AX(state=\texttt{DISPATCHING}))$
V Property Specification
The formalization of the HA and Task Lifecycle models provides the foundation for analyzing and verifying critical system properties. In this context, formal properties are verifiable statements about the system’s execution paths, used to ensure the safety, security, and functionality of an Agentic AI system’s behavior. We structure these properties into two distinct sets: those captured by the HA model (overall orchestration, Table I), and those governing the task lifecycle model (sub-task state transitions, Table II).
These properties do not operate in isolation; they form a tightly coupled system of interdependent guarantees necessary for correct and reliable function in a multi-AI agent environment. Both sets of properties are expressed using temporal logics, Computation Tree Logic (CTL) and Linear Temporal Logic (LTL). These logics allow for the precise, unambiguous definition of desired system behavior over time and execution paths.
V-A Property Design Principles and Definitions
Both models include a set of property categories, each containing multiple specific properties, to establish verifiable guarantees. The core categories used in both tables are liveness, safety, ordering, completeness, fairness, and reachability. These properties ensure the overall system makes progress, avoids undesirable states, achieves its goals, and handles reliable resource allocation.
Liveness. Liveness ensures that an Agentic AI system will eventually make progress toward a goal. This property guarantees that, despite asynchronous actions, inter-agent dependencies, or potential conflicts, the system avoids global deadlock or infinite starvation. This is critical in these systems, where a lack of progress in one agent can propagate and result in a cascading stall of the entire system.
  • Example (Host Agent Model, $\mathtt{HP_{1}}$): Every user prompt ($Req_{U}$) must eventually receive a final response ($Resp_{H}$) from the host agent. This is expressed in temporal logic as $AG(Req_{U}\to AF\ Resp_{H})$.
  • Example (Task Lifecycle Model, $\mathtt{TL_{1}}$): A sub-task that is CREATED must eventually terminate in a defined end state (COMPLETED, ERROR, or CANCELED).
Safety. Safety ensures that agents never enter globally invalid or harmful states, even when operating asynchronously or under adversarial conditions. In an Agentic AI system, this includes verifying that an agent’s actions do not cause irreversible policy violations in the collective system state.
To formally reason about the safety of the system, we introduce the Validation Module (VM). The VM is responsible for enforcing trust and validation constraints on external entities (EEs). We define $VM(\text{EE})$ as the proposition that a given entity EE has successfully passed all predefined validation processes, including schema validation and behavioral assessments. For the formal analysis, we assume the existence of a correct and functioning VM.
  • Example (Host Agent Model, $\mathtt{HP{{}_{9}}}$): This property ensures task invocation ($\mathsf{CL.invoke}$) is strictly conditional on the External Entity (EE) having completed system validation ($VM(\text{EE})$). Formally, $AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$.
[…]
Ordering (Mutual Exclusion). This property ensures agents avoid simultaneous access to exclusive resources and execute dependent actions in the correct sequence. The Orchestrator must strictly enforce the execution order of sub-tasks across specialized AI agents according to the DAG.
  • Example (Host Agent Model, $\mathtt{HP_{10}}$): A sub-task is invoked only when it has no uncompleted dependencies, preventing the consumption of intermediate data that is still being updated. This is expressed as $\bigwedge_{i=1}^{n=|D|}AG(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{sub\_task}_{i}))\to\text{dependencies}[\text{sub\_task}_{i}]=\emptyset$.
Completeness. Completeness captures the guarantee that if a valid solution or coordinated plan exists, the multi-AI agent system will find it. This property is relevant for distributed planning, where incomplete reasoning could leave solvable tasks unfulfilled due to suboptimal agent coordination.
  • Example (Host Agent Model, $\mathtt{HP{{}_{1}2}}$): Every user prompt ($Req_{U}$) must lead either to intent clarification ($\mathsf{Clarify\_Intent}$) or to task planning ($\mathsf{LLM.Build\_Task\_DAG}$), ensuring the system processes every request and initiates a path toward a solution. This behavior is captured as $AG(Req_{U}\to EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\vee\mathsf{Clarify\_Intent}))$.
Fairness. Fairness requires that all sub-tasks delegated to agents must eventually transition to a completed or error state, or be explicitly terminated (e.g., via a timeout). This prevents a single agent’s failure from causing a system-wide deadlock, ensuring the overall task can progress.
[…]
Reachability. Reachability requires that desired joint states or configurations are achievable through the agents’ available actions and communication patterns. In multi-AI agent coordination, this involves verifying that the combination of capabilities and allowed transitions can lead from the current distributed state to the intended goal without deadlocks or inescapable loops.
  • Example (Host Agent Model, $\mathtt{HP_{16}})$: It is always possible to reach a state where the host agent replies to the user. This property is expressed as $EF(Resp_{H})$.
VI Case Study: Verification of Architectural Controls against Adversarial Behavior
The host agent and task lifecycle models provide a framework to detect, constrain, and mitigate adversarial behaviors in Agentic AI systems. Here, adversarial behavior is defined as any external deviation that compromises correctness, violates system invariants, or exploits control flow inconsistencies.
These models can be leveraged to employ a layered security architecture, where security constraints are encoded as temporal logic invariants at distinct architectural layers. This structure enables verification of system-level defenses, including secure communication, trust anchoring, intent integrity, and failure containment.
Control Point 1: Host Agent Core (Intent Integrity) Securing the primary human-AI interface, the HAC, serves as the initial control point. The core architecture is vulnerable to attacks such as prompt injection [39, 40, 14] and jailbreaks [41] that subvert intent integrity.
Our model establishes a security boundary by leveraging the explicit “Clarify Intent” phase as a guard. The verification process focuses on ensuring all requests lead to a traceable outcome and enforces intent authenticity.
  • Verifying Correctness (Completeness, $\mathtt{HP_{12}}$): Verifies that every user prompt initiates a clear path toward a solution (clarification or planning), which prevents the core from silently discarding or misinterpreting a request. This property is defined as $AG(Req_{U}\rightarrow EX(\mathsf{LLM.Build\_Task\_DAG}(\mathcal{I}_{U},\{\mathcal{EE}_{info}\}))\lor\mathsf{Clarify\_Intent}))$.
  • Verifying Progress (Liveness, $\mathtt{{HP_{2}}}$): Guarantees the HAC eventually clarifies the intent of each user prompt ($AG(Req_{U}\rightarrow AF\>\mathcal{I}_{U})$), and prevents denial-of-service via indefinite ambiguity.
Control Point 2: Registry (Trust Anchoring) The Registry functions as the trust anchor for all EE interactions. This mitigates the supply chain risk of integrating malicious components [42, 43]. Our model ensures the integrity of service discovery and registration via codified security properties:
  • Verifying Trust (Safety, $\mathtt{{HP_{9}}}$): Enforces that task invocation ($\mathsf{CL.invoke}$) is strictly conditional on the EE completing system validation ($VM(\text{EE})$). This formal verification ensures runtime decisions adhere to established trust requirements, preventing privilege escalation by unvetted entities. Formally: $AG\>(\mathsf{CL.invoke}(\text{EE},\text{protocol},\text{payload})\rightarrow VM(\text{EE}))$.
Control Point 3: Orchestrator (Delegation Monitoring) The Orchestrator serves as a runtime security monitor to enforce verifiable security boundaries during task execution through the dependency DAG structure. This layer directly defends against coordination-based threats.
  • Verifying Integrity and Ordering (Safety, $\mathtt{HP_{10}}$): The DAG structure guarantees strict ordering. The property $\mathtt{HP_{10}}$ ensures task invocation occurs only when all uncompleted dependencies are resolved, and prevents premature execution and data inconsistency.
  • Verifying Causal Isolation (Failure Containment): The Orchestrator’s dependency management enforces strict causal isolation. This confines adversarial effects within the minimal execution subgraph (e.g., a sub-task will not proceed if any dependency state is FAILED or INVALIDATED). This guarantees fault propagation is contained, and supports system liveness and safety.
Control Point 4: Communication Layer and Zero-Trust The CL acts as the secure unified protocol security infrastructure, and enforces a zero-trust model for all protocol-level exchanges (REST, gRPC, etc.).
  • Formal Requirements: The $\mathsf{CL.invoke}()$ operation with target, protocol, and payload arguments must satisfy non-negotiable security policies to ensure that any received data or instruction is authentic and trustworthy. These policies include verifiable identity, message integrity, and confidentiality.
  • Verifying Reliability (Fairness, $\mathtt{HP_{13}}$, $\mathtt{HP_{14}}$): The fairness properties implicitly verify the reliability of this infrastructure by requiring that all A2A agent RPC calls and MCP JSON-RPC calls eventually receive responses, ensuring continuous, reliable operation of the communication channel.
VII Related Work
Multi-Agent System Architectures and Formal Models. Traditional Multi-Agent Systems (MAS) research established foundational principles for distributed autonomous systems, focusing on authorization and secure communication between agents [16]. The development of Large Agentic AI systems has introduced new architectural challenges, necessitating specific coordination frameworks. Surveys have provided valuable taxonomies of agent systems [44, 45, 26]; however, these works largely describe system categories rather than offering operational models that enable property specification, formal reasoning, or verification. Our approach builds upon existing architectural insights but advances them by proposing the models for the host agent and task lifecycle.
The Challenge of Integrated Protocol Security. The safety and security of Agentic AI systems is a major research domain that directly motivates our formal framework. Security challenges in these systems have been comprehensively documented, identifying vulnerability classes such as coordination attacks, information leakage, and privilege escalation through delegation chains [6]. This work highlights the cascade failure phenomenon, where the compromise of a single agent propagates across the entire system. Industry assessments, such as the OWASP Foundation’s draft “Top 10 for Agentic AI”, codified these concerns and cite “Unreliable Delegation & Coordination” as a critical risk that directly relates to the heterogeneous protocol integration problem we address [46].
Protocol-Specific Vulnerabilities. The need for formal verification is compounded by the rapid emergence of vulnerabilities in key communication protocols. Research on the MCP identified major threats, including installer spoofing, sandbox escape, and critical gaps in privilege management in real-world applications [42, 47]. Furthermore, studies demonstrate that LLMs can be coerced into executing malicious operations via MCP tools, giving rise to retrieval-agent deception attacks and systematic security benchmarks [48, 49]. Similarly, foundational security analysis of the A2A Protocol uncovered vulnerabilities in identity management and task exchange protocols [19]. Key risks in these multi-agent communications include ensuring message integrity and agent authentication, with proposed improvements focusing on validating agent identities and protecting data exchange to mitigate manipulation [21]. A critical threat is prompt infection [14], where malicious instructions propagate between AI agents, creating persistence across tasks. This can enable a single compromised log file to trigger a cascading attack, potentially escalating into a company-wide ransomware incident. Empirical studies confirm that malicious agents can exploit delegation mechanisms and use coordinated manipulation to subvert system goals, underscoring the insufficiency of traditional security testing for cross-protocol interactions [12, 17, 10]. This body of work confirms the urgent need for a unified semantic framework capable of verifying integrated A2A and MCP interactions.
Formal Verification and Property Specification. Our methodology employs property specification, which is validated by a history of research in autonomous systems. Early work established security models for distributed AI systems that focus on the foundational security principles of authentication and authorization [16]. More recently, the focus shifted to the rigorous application of formal methods to verify system correctness and reliability. The importance of this approach for agent coordination protocols is widely recognized [46]. Our use of formal logic and property categorization (e.g., liveness, safety) provides the rigorous foundation necessary to prove end-to-end correctness, distinguishing our work from empirical studies that demonstrate failure without offering a mathematically verifiable solution.
VIII Conclusion
The proliferation of communication and coordination protocols in Agentic AI systems creates a semantic gap that prevents the rigorous analysis of system safety, security, and functionality. This paper addresses that gap by introducing a modeling framework that consists of two foundational models: the host agent, which orchestrates user-prompted tasks, and the task lifecycle, which details sub-task execution. Grounded in this framework, we define $\mathbf{31}$ formal properties, spanning liveness, safety, completeness, and fairness, expressed in temporal logic. These properties enable the formal verification of system behavior, detection of coordination edge cases, and prevention of deadlocks. Our work provides a domain-agnostic approach for designing and deploying verifiably safe and reliable Agentic AI systems. To operationalize this methodology, our future work will focus on automatically detecting property violations in coded Agentic AI systems by deriving a formal model from the code and model-checking specified properties against it.